# Unifying Demonstration Selection and Compression for In-Context Learning

### v2

@@ 4.1 Baselines
Naive Vicuna-7b serves as the fundamental baseline fed into actual demonstrations. AutoCompressor recurrently compresses demonstrations into virtual tokens. We employ their Llama2-7b version [^5] https://github.com/princeton-nlp/AutoCompressors.. LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words. We employ their released 7b version [^6] https://github.com/microsoft/LLMLingua., of which compressor is a fine-tuned Llama-2. For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b. ICAE [^7] https://github.com/getao/icae. compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b.
Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
@@ 4.2 Settings
Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 512 for both compression and generation. For a fair comparison, we set the allowed window of baselines to 512 and their compression ratio to the same as UniICL. We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation). Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration because of marginal ICL gains and memory costs.
We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input for all baselines. UniICL is employed to perform selection among them in practice due to computation efficiency for high-resource ICL. On the contrary, the low-resource ICL setting fixes the random candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
@@ 4.3 Results
Model #-shots CoLA-dev SST-2-dev IMDb ARXIV XSum Acc. R-1 R-2 R-L R-1 R-2 R-L Vicuna 0-shot 56.2 91.7 92.6 34.3 9.1 27.4 19.9 5.0 13.4 1-shot 58.2 (57.4) 90.7 (90.8) 91.9 (91.0) 34.4 (33.2) 9.1 (8.5) 27.5 (26.7) 21.2 (20.4) 5.8 (5.2) 14.5 (13.9) 2-shot 62.1 (59.8) 92.1 (91.3) 91.7 (91.7) - - - - - - 5-shot 62.3 (61.9) 93.0 (91.9) 94.1 (92.5) - - - - - - AutoCompressor 1-shot 42.1 (40.9) 85.7 (84.2) 95.0 (95.1) 27.0 (26.4) 8.4 (8.2) 26.1 (25.8) 21.3 (20.3) 6.5 (6.3) 13.7 (13.7) 2-shot 58.8 (56.3) 88.0 (86.4) 95.0 (94.6) 27.1 (26.2) 8.6 (7.9) 26.4 (25.4) 21.9 (21.4) 6.6 (6.4) 14.5 (14.1) 5-shot 59.1 (58.8) 91.3 (89.1) 94.7 (94.8) 34.5 (33.7) 9.4 (9.1) 28.7 (27.9) 22.4 (21.7) 6.9 (6.7) 14.8 (14.3) LLMLingua 1-shot 55.5 (55.0) 89.7 (89.6) 91.0 (89.9) 33.3 (33.1) 8.9 (8.7) 27.4 (27.1) 20.5 (19.7) 5.4 (5.2) 14.5 (14.4) 2-shot 56.7 (55.7) 90.7 (90.2) 91.3 (91.0) 32.9 (32.0) 8.2 (8.1) 26.9 (25.9) 20.3 (20.0) 5.2 (5.1) 14.3 (14.1) 5-shot 57.2 (56.9) 90.6 (90.2) 90.9 (91.2) 30.1 (29.7) 7.9 (7.4) 25.3 (24.6) 19.7 (18.6) 4.9 (4.9) 14.1 (14.3) ICAE 1-shot 30.9 (30.9) 61.0 (60.1) 85.7 (83.3) 26.8 (24.6) 8.2 (7.1) 24.7 (22.9) 23.5 (21.9) 8.5 (7.8) 20.9 (20.3) 2-shot 30.9 (30.9) 49.0 (52.8) 85.9 (85.9) 27.2 (25.5) 8.4 (7.6) 25.9 (24.3) 24.4 (23.2) 8.9 (8.4) 21.3 (20.8) 5-shot 30.9 (30.9) 54.2 (51.0) 85.7 (85.9) 28.3 (26.9) 8.7 (7.7) 26.6 (25.8) 25.3 (24.9) 9.2 (8.8) 22.5 (21.6) UniICL 1-shot 58.7 (58.0) 92.9 (91.7) 94.3 (92.3) 35.5 (34.7) 10.5 (10.2) 28.7 (27.9) 27.7 (25.5) 10.2 (9.1) 21.2 (20.0) 2-shot 62.4 (61.0) 92.4 (91.6) 94.9 (93.3) 36.1 (35.2) 10.8 (10.4) 29.4 (28.2) 29.4 (26.8) 11.0 (9.8) 22.3 (20.9) 5-shot 62.6 (61.8) 93.1 (92.3) 94.5 (94.0) 35.8 (35.4) 10.6 (10.2) 29.5 (28.1) 30.7 (27.6) 11.3 (10.1) 22.8 (21.4) UniICL 1-shot 59.1 (58.7) 93.0 (91.9) 94.5 (91.6) 34.8 (34.7) 10.4 (10.3) 28.1 (27.8) 29.1 (26.2) 10.8 (9.4) 22.2 (20.7) 2-shot 62.6 (61.2) 94.0 (93.0) 94.9 (92.3) 34.6 (34.3) 10.6 (10.4) 28.5 (28.3) 30.3 (28.9) 11.3 (10.5) 22.9 (21.7) 5-shot 63.3 (61.5) 94.7 (92.8) 95.0 (93.8) 35.6 (35.3) 11.0 (10.8) 29.1 (27.7) 31.1 (30.0) 11.7 (11.2) 23.5 (22.3) 8-shot 63.8 (62.6) 94.7 (93.1) 95.0 (94.2) - - - - - - UniICL + $L_{ctr}$ 1-shot 59.3 (58.9) 93.2 (92.4) 95.1 (92.8) 35.6 (35.1) 10.7 (10.5) 28.9 (28.3) 30.0 (27.9) 11.3 (10.1) 22.8 (21.5) 2-shot 62.4 (62.0) 94.5 (92.8) 94.8 (93.4) 36.8 (35.3) 10.8 (10.6) 29.6 (28.9) 30.8 (29.2) 11.4 (10.7) 23.0 (21.9) 5-shot 64.3 (61.8) 94.7 (93.4) 96.1 (94.2) 37.1 (34.9) 11.3 (11.2) 30.0 (29.3) 32.5 (30.6) 12.3 (11.8) 24.7 (23.8) 8-shot 64.7 (63.3) 94.7 (94.1) 95.6 (95.0) - - - - - -
@@
Table 2: The high- and low-ICL results on CoLA-dev, SST-2-dev, and IMDb. Results in () represent low-resource ICL. represents the demonstrations selected by UniICL, and the others are selected by S-BERT. +$L_{ctr}$ indicates the selection augmented UniICL (optimized with Equation 7). Bold (underline) represents the best performance on high- and low-resource ICL.
@@
We comprehensively evaluate the ICL performance of UniICL on the out-domain dataset CoLA, SST-2, and IMDb by close-ended evaluation and ARXIV by open-ended evaluation, as demonstrated in Table 2. Specifically, UniICL outperforms naive Vicuna-7b fed with actual candidate demonstrations, which indicates that virtual demonstrations are more efficient and informative for guiding the target LLM, and UniICL outperforms all the baselines by compressing the same demonstrations pre-ranked by S-BERT. Additionally, UniICL achieves further performance gains after selecting demonstrations via itself. Additionally, open-ended results indicate that virtual demonstrations still efficiently capture semantic information for ICL guiding, even though summarization demonstrations are much longer than understanding ones. Regarding ARXIV, the original ICL is not helpful enough due to its extremely over-length document, leaving little room for demonstrations. UniICL works as expected by compressing demonstrations and concatenating virtual demonstrations, and achieves +2.8 R-1 gains in the 5-shot setting that selects demonstrations via selection-augmented virtual tokens.
Furthermore, The ablation experiments of +$\mathcal{L}_{ctr}$ show that UniICL is faced with performance degradation without $L_{ctr}$ and the performance gap becomes larger with the number of demonstrations increasing. The results of BlueLM are exhibited in Appendiex B.
@@
Method MS MARCO
BM25 18.5
Vicuna 28.9
AutoCompressor 29.3
ICAE 30.2
UniICL 31.6
Table 3: Results on MS MARCO. Vicuna applies the last hidden states of [EOS] to represent sentences in latent space. Following the previous study, we report MRR@10. $\dagger$ means citing from Liang Wang et al. (2022).
Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-domain MS MARCO dataset in Table 3. UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods. Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al. (2024) since most of them are fine-tuned on this dataset, which is unfair for comparison.
Figure 7: The overall sensitivity analysis of compression ratio.
@@ 5.1 Compression Ratio
During training, the compression ratio is dynamically sampled from 2 to 16. We mix up 2,000 instances from the in-domain validation set, 1,000 for XSUM, and 1,000 for CICERO to select the compression ratio for UniICL in Figure 7, with the backbone of Vicuna and BlueLM respectively. Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed. Therefore, we can measure the dense information quality of the same content with different compression ratios by ROUGE-1 since it is more sensitive to token-level differences. The performance is relative smoothing when the compression ratio changes from 4×4$\times$4 × to 12×12$\times$12 ×. However, when it comes to 16×16$\times$16 ×, an obvious drop occurs. Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments. The 512×512$\times$512 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 512.
Figure 8: The efficiency comparison between UniICL and other compression methods in CoLA with the number of shots increasing from 0 to 64. Memory explodes are represented as *, corresponding to the break of the line chart.
@@ 5.2 Efficiency Analysis
In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 0.24%. We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 8. With the help of the Demonstration Bank (DB), UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching). Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline). What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
Additionally, We demonstrate the inference computation and GPU hours in Table 4, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens. Notably, UniICL (without DB) compresses the former half, and the latter half is fed into the generator directly, while Vicuna and Vicuna-1k are distinguished in window limitations. Results indicate that minimal GPU hours increased due to the parallel computation of forward, although the extra compression of UniICL surges the computation. Additionally, Vicuna with a 1k window limitation surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.
5.3 Training Deviation
To quantify the performance gains brought by the learnable projection layer. We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA in Table 5, setting the rank to 32. UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.

### v3

@@ 4.1 Baselines
Unmodified Vicuna-7b serves as the fundamental baseline fed with actual demonstrations. AutoCompressor compresses prompts into 50 virtual tokens in different rounds recurrently. Previous compressed virtual tokens are put at the beginning of the current segment. Finally, virtual tokens of different compression rounds are concatenated for generation. We employ their Llama2-7b version for comparison. LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words. We employ their released 7b version, of which the compressor is a fine-tuned Llama2. For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b. ICAE compresses demonstrations into 128 virtual tokens via a LoRA-adapted Llama2-7b. Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
@@ 4.2 Settings
Dataset # words (96,512] (512,1024] (1024,1536] XSum Narayan et al. (2018) - 10,000 4,697 CICERO Ghosal et al. (2022) 10,000 - - SUPER-NI Wang et al. (2022b) - 10,000 7,000 XSum (Ctr) 5,000
Table 1: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSum (Ctr) is used for the second-phase training in Eq. 6.
Dataset In-Domain # Test # Demonstrations MS MARCO-dev 6,980 - XSum 1,500 204,045/20 Arxiv 1,500 203,037/20 CoLA-dev 1,041 67,349/20 SST-2-dev 872 8,551/20 IMDb 1,500 25,000/20 MMLU 13,985 25,000/20
Table 2: The details of the involved evaluation datasets. -dev represents employing the development set due to their test sets are inaccessible. # Demonstrations represent the number of demonstrations to be selected in high/low-resource ICL settings.
We construct the training set by mixing up XSum, CICERO, and SUPER-NI according to their length as shown in Tab. 1 and evaluate UniICL on extensive out-of-domain datasets as listed in Tab. 2, with more details reported in Appendix H. Considering computation efficiency, we set the max allowed input length limit to 512 for both compression and generation for both training and inference. For a fair comparison, we set the allowed window of baselines to 512, and the compression ratio of default UniICL and baselines is set to 12, which is determined by the validation in Fig. 7. We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation). We train 10 epochs and 2 epochs respectively for the first- and second-phase training. The best checkpoints are selected according to their performance on in-domain validation sets. Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shots to 8 for understanding tasks and 5 for generative tasks for illustration, because of marginal ICL gains and memory costs.
We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input for all baselines. UniICL is employed to perform selection among them in practice due to computational efficiency for high-resource ICL. On the contrary, the low-resource ICL setting utilizes the randomly sampled 20 candidate demonstrations for all inference inputs, while UniICL performs selection as normal.
To verify the universality, we further build UniICL on BlueLM-7B Team (2023) and Llama2-7B Touvron et al. (2023). Results of BlueLM and Llama2 will be reported in Appendix C and Appendix D.
@@ 4.3 Results
Model #-shots CoLA-dev SST-2-dev IMDb Arxiv XSum Acc. R-1 R-2 R-L R-1 R-2 R-L Vicuna 0-shot 56.2 91.7 92.6 34.3 9.1 27.4 19.9 5.0 13.4 1-shot 58.2 (57.4) 90.7 (90.8) 91.9 (91.0) 34.8 (34.4) 9.3 (9.1) 27.9 (27.5) 21.5 (21.2) 5.9 (5.8) 14.7 (14.5) 2-shot 62.1 (59.8) 92.1 (91.3) 91.7 (91.7) - - - - - - 5-shot 62.3 (61.9) 93.0 (91.9) 94.1 (92.5) - - - - - - AutoCompressor 1-shot 42.1 (40.9) 85.7 (84.2) 95.0 (95.1) 27.0 (26.4) 8.4 (8.2) 26.1 (25.8) 21.3 (20.3) 6.5 (6.3) 13.7 (13.7) 2-shot 58.8 (56.3) 88.0 (86.4) 95.0 (94.6) 27.1 (26.2) 8.6 (7.9) 26.4 (25.4) 21.9 (21.4) 6.6 (6.4) 14.5 (14.1) 5-shot 59.1 (58.8) 91.3 (89.1) 94.7 (94.8) 34.5 (33.7) 9.4 (9.1) 28.7 (27.9) 22.4 (21.7) 6.9 (6.7) 14.8 (14.3) LLMLingua 1-shot 55.5 (55.0) 89.7 (89.6) 91.0 (89.9) 33.3 (33.1) 8.9 (8.7) 27.4 (27.1) 20.5 (19.7) 5.4 (5.2) 14.5 (14.4) 2-shot 56.7 (55.7) 90.7 (90.2) 91.3 (91.0) 32.9 (32.0) 8.2 (8.1) 26.9 (25.9) 20.3 (20.0) 5.2 (5.1) 14.3 (14.1) 5-shot 57.2 (56.9) 90.6 (90.2) 90.9 (91.2) 30.1 (29.7) 7.9 (7.4) 25.3 (24.6) 19.7 (18.6) 4.9 (4.9) 14.1 (14.3) ICAE 1-shot 30.9 (30.9) 61.0 (60.1) 85.7 (83.3) 26.8 (24.6) 8.2 (7.1) 24.7 (22.9) 23.5 (21.9) 8.5 (7.8) 20.9 (20.3) 2-shot 30.9 (30.9) 49.0 (52.8) 85.9 (85.9) 27.2 (25.5) 8.4 (7.6) 25.9 (24.3) 24.4 (23.2) 8.9 (8.4) 21.3 (20.8) 5-shot 30.9 (30.9) 54.2 (51.0) 85.7 (85.9) 28.3 (26.9) 8.7 (7.7) 26.6 (25.8) 25.3 (24.9) 9.2 (8.8) 22.5 (21.6) UniICL 1-shot 58.7 (58.0) 92.9 (91.7) 94.3 (92.3) 35.5 (34.7) 10.5 (10.2) 28.7 (27.9) 27.7 (25.5) 10.2 (9.1) 21.2 (20.0) 2-shot 62.4 (61.0) 92.4 (91.6) 94.9 (93.3) 36.1 (35.2) 10.8 (10.4) 29.4 (28.2) 29.4 (26.8) 11.0 (9.8) 22.3 (20.9) 5-shot 62.6 (61.8) 93.1 (92.3) 94.5 (94.0) 35.8 (35.4) 10.6 (10.2) 29.5 (28.1) 30.7 (27.6) 11.3 (10.1) 22.8 (21.4) UniICL 1-shot 59.1 (58.7) 93.0 (91.9) 94.5 (91.6) 34.8 (34.7) 10.4 (10.3) 28.1 (27.8) 29.1 (26.2) 10.8 (9.4) 22.2 (20.7) 2-shot 62.6 (61.2) 94.0 (93.0) 94.9 (92.3) 34.6 (34.3) 10.6 (10.4) 28.5 (28.3) 30.3 (28.9) 11.3 (10.5) 22.9 (21.7) 5-shot 63.3 (61.5) 94.7 (92.8) 95.0 (93.8) 35.6 (35.3) 11.0 (10.8) 29.1 (27.7) 31.1 (30.0) 11.7 (11.2) 23.5 (22.3) 8-shot 63.8 (62.6) 94.7 (93.1) 95.0 (94.2) - - - - - - UniICL + $L_{ctr}$ 1-shot 59.3 (58.9) 93.2 (92.4) 95.1 (92.8) 35.6 (35.1) 10.7 (10.5) 28.9 (28.3) 30.0 (27.9) 11.3 (10.1) 22.8 (21.5) 2-shot 62.4 (62.0) 94.5 (92.8) 94.8 (93.4) 36.8 (35.3) 10.8 (10.6) 29.6 (28.9) 30.8 (29.2) 11.4 (10.7) 23.0 (21.9) 5-shot 64.3 (61.8) 94.7 (93.4) 96.1 (94.2) 37.1 (34.9) 11.3 (11.2) 30.0 (29.3) 32.5 (30.6) 12.3 (11.8) 24.7 (23.8) 8-shot 64.7 (63.3) 94.7 (94.1) 95.6 (95.0) - - - - - -
@@
Table 3: The high- and low-ICL results on CoLA-dev, SST-2-dev, and IMDb. Results in (bracket) represent low-resource ICL. represents the demonstrations selected by UniICL, and the others are selected by S-BERT. +$L_{ctr}$ indicates the selection augmented UniICL (optimized with Eq. 6). Bold (underline) represents the best performance on high- and low-resource ICL. R- indicates Rouge scores. All compression methods are evaluated with a compression ratio set to 12.
@@
We comprehensively evaluate the ICL performance of UniICL on the out-of-domain dataset CoLA, SST-2, and IMDb by close-ended evaluation and Arxiv by open-ended evaluation in Tab. 3. The details of the involved evaluation datasets and metrics are reported in Tab. 2 and Appendix H. Specifically, UniICL outperforms unmodified Vicuna-7b fed with actual candidate demonstrations, which indicates that Memory Tokens are more efficient and informative for guiding the target LLM. Meanwhile, UniICL outperforms all the baselines by compressing the same demonstrations pre-ranked by S-BERT. Additionally, UniICL achieves further performance gains after selecting demonstrations via itself (UniICL). The open-ended results highlight that Memory Tokens indeed capture semantic information for ICL generation, even though summarization demonstrations are much longer than understanding ones. Regarding Arxiv, the original ICL is not helpful enough due to its extremely over-length document, leaving little room for demonstrations. UniICL works as expected by compressing demonstrations into Memory Tokens and concatenating them, achieving +2.8 Rouge-1 gains in selection-augmented UniICL (+$\mathcal{L}_{ctr}$). Additionally, according to the results of +$\mathcal{L}_{ctr}$, we find that the gains brought by selection augmentation become larger as the number of demonstrations increases. We attribute this to the fact that UniICL selects more useful demonstrations for generation after the second-phase training. The results of BlueLM are exhibited in Appendix C. Except for understanding and generative tasks, we further evaluate UniICL on MMLU in Tab. 4. UniICL achieves stable performance gains with more demonstrations introduced. Additionally, considering ICAE and AutoCompressor are soft-prompt-based compression methods built on Llama2, we also build UniICL on Llama2 for ablation in Appendix D.
#-Shots S H SS O Avg. 0-shot 36.9 53.2 53.7 50.7 48.6 1-shot 38.6 55.3 54.6 52.4 50.2 2-shot 39.2 55.8 55.3 53.1 50.9 5-shot 40.1 55.6 55.3 53.8 51.2
Table 4: Performance of UniICL on MMLU benchmark. We reported the Accuracy at the category level. S represents STEM, H represents Humanities, SS represents Social Science, O represents Other, and Avg indicates their average performance.
@@
Method # TP MRR@10 BM25 - 18.5 Vicuna - 28.9 AutoCompressor - 29.3 ICAE - 30.2 UniICL - 31.6 SIMLM†‡ 110M 41.1 UniICL 17M 38.9
Table 5: MRR@10 results on MS MARCO. Vicuna applies the last hidden states of [EOS] to represent sentences in latent space. Results citing from Liang Wang et al. (2022a) are denoted as , and methods supervised trained on MS MARCO are represented as . Bold indicates the best zero-shot performance and Underline is the best fine-tuned results. # TP indicates the number of trainable parameters.
Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-of-domain MS MARCO dataset in Tab. 5. UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods. Subsequently, we fine-tune the first-phase compression model of UniICL on the training set of MS MARCO. UniICL achieves comparable performance with SIMLM Wang et al. (2022a), which is specified in Information Retrieval (IR) and has more trainable parameters.
Figure 7: The compression ratio sensitivity analysis of Llama2 , BlueLM, and Vicuna.
@@ 5.1 Compression Ratio
#-shots CoLA SST-2 IMDb Arxiv Acc. R-1 1-shot 58.5 (-0.8) 91.4 (-1.8) 92.6 (-2.5) 34.8 (-0.8) 2-shot 59.7 (-2.7) 92.1 (-2.4) 94.1 (-0.7) 35.7 (-1.1) 5-shot 62.4 (-1.9) 93.1 (-1.6) 94.8 (-1.3) 36.6 (-0.5)
Table 6: Performance of UniICL on out-of-domain datasets, with a fixed compression ratio set to 12 during training.
During training, the compression ratio is dynamically sampled from 2 to 16. We mix up 2,000 instances from the in-domain validation set, 1,000 for XSum, and 1,000 for CICERO to select the compression ratio for UniICL in Fig. 7, with the backbone of Llama2, Vicuna, and BlueLM respectively. Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed. Therefore, we can measure the dense information quality of the same content with different compression ratios by ROUGE-1 since it is more sensitive to token-level differences. The performance is relative smoothing when the compression ratio changes from 4×4$\times$4 × to 12×12$\times$12 ×. However, when it comes to 16×16$\times$16 ×, an obvious drop occurs. In order to analyze this phenomenon more deeply, we provide a thorough analysis in Appendix G. Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments. The 512×512$\times$512 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 512.
Figure 8: The efficiency comparison between UniICL and other compression methods in CoLA with the number of shots increasing from 0 to 64. Memory explodes are represented as *, corresponding to the break of the line chart. +Caching represents using DB.
To explore whether it could yield additional performance gains compared with dynamic ratios, in Tab. 6, we re-train UniICL with the compression ratio fixed to 12 (Results of more fixed ratios are reported in Appendix F.). Results indicate that UniICL trained with fixed compression ratios underperforms in out-of-domain datasets as it exhibits over-fitting in in-domain sets as shown in Tab. 11.
Furthermore, we analyze whether 12$\times$ is suitable for all out-of-domain datasets in Fig. 9 in Appendix E. Results indicate that 12$\times$ outperforms other compression ratios in general across 4 out-of-domain datasets. It also points out that lower ratios still work comparable for short demonstrations and higher ratios are suitable for long demonstrations to some extent.
@@ 5.2 Efficiency Analysis
In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 0.24%. We evaluate the memory costs and inference latency of UniICL and other compression methods in Fig. 8. With the help of the Demonstration Bank (DB), UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching). Despite this, parallel computation facilitates the compression process, resulting in minimal throughput degradation (UniICL and Baseline). The unmodified 7B LLM causes a memory explosion for 8-shot settings, and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shot within a 24GB CUDA allocation.
Additionally, we demonstrate the inference computation and GPU hours in Tab. 7, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens. Notably, UniICL (without DB) compresses the former half, and the latter half is fed into the generator directly, while Vicuna and Vicuna-1k are distinguished in window limitations. Results indicate that minimal GPU hours increased due to the parallel computation of forward, although the extra compression of UniICL surges the computation. Additionally, Vicuna, with a 1k window limitation, surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.